English

The odds against student X solving a business statistics problem are 8 : 6 and odds in favour of student Y solving the same problem are 14 : 16 What is the probability that neither solves the problem?

Advertisements
Advertisements

Question

The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the probability that neither solves the problem?

Diagram
Sum
Advertisements

Solution

Let A be the event that X solves the problem B be the event that Y solves the problem.
Since the odds against student X solving the problem are 8: 6
∴ Probability of occurrence of event A is given by.

P(A) = `6/(8 + 6) = 6/14` and

P(A') = 1 – P(A) = `1 - 6/14 = 8/14`

Also, the odds in favour of student Y solving the problem are 14: 16
∴ Probability of occurrence of event B is given by

P(B) = `14/(14 + 16) = 14/30` and

P(B') = 1 – P(B) = `1 - 14/30 = 16/30`

Now A and B are independent events.
∴ A' and B' are independent events.

P (neither solves the problem) = P(A' ∩ B')
= P(A') . P(B')

= `8/14 xx 16/30`

= `32/105`

shaalaa.com
  Is there an error in this question or solution?
Chapter 7: Probability - Miscellaneous Exercise 7 [Page 110]

APPEARS IN

Balbharati Mathematics and Statistics (Commerce) Part 2 [English] Standard 11 Maharashtra State Board
Chapter 7 Probability
Miscellaneous Exercise 7 | Q 14. (b) | Page 110

RELATED QUESTIONS

Let E and F be events with `P(E) = 3/5, P(F) = 3/10 and P(E ∩ F) = 1/5`.  Are E and F independent?


Bag A contains 3 red and 2 white balls and bag B contains 2 red and 5 white balls. A bag is selected at random, a ball is drawn and put into the other bag, and then a ball is drawn from that bag. Find the probability that both the balls drawn are of same color


Solve the following:

Let A and B be independent events with P(A) = `1/4`, and P(A ∪ B) = 2P(B) – P(A). Find `"P"("B'"/"A")`


If A and B are independent events such that P(A) = p, P(B) = 2p and P(Exactly one of A, B) = `5/9`, then p = ______.


Three events A, B and C are said to be independent if P(A ∩ B ∩ C) = P(A) P(B) P(C).


Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: 1 – (1 – P1)(1 – P2) 


If A and B are two events and A ≠ Φ, B ≠ Φ, then ______.


If A and B are independent events, then A′ and B′ are also independent


If A and B are two independent events then P(A and B) = P(A).P(B).


Let A and B be two events. If P(A | B) = P(A), then A is ______ of B.


The probability that A hits the target is `1/3` and the probability that B hits it, is `2/5`. If both try to hit the target independently, find the probability that the target is hit.


Given two independent events, if the probability that exactly one of them occurs is `26/49` and the probability that none of them occurs is `15/49`, then the probability of more probable of the two events is ______.


Given two events A and B such that (A/B) = 0.25 and P(A ∩ B) = 0.12. The value P(A ∩ B') is ______.


Two events E and F are independent when which conditional probability condition holds, provided \[P(E) \ne 0\]?


Two events E and F are independent when which conditional probability condition holds, provided \[P(F) \ne 0\]?


A die is thrown once. If E is the event that the number appearing is a multiple of 3, which set represents E?


A die is thrown once. If F is the event that the number appearing is even, which set represents F?


For \[E=\{3,6\}\] when a die is thrown once, what is \[P(E)\]?


For \[E\cap F=\{6\}\] when a die is thrown once, what is \[P(E\cap F)\]?


Given \[P(E\cap F)=\frac{1}{6}\] and \[P(E)P(F)=\frac{1}{6}\], what is the relation between E and F?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×